40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jan 26 17:50:22 2016, gautam, Update, Green Locking, Lightwave frequency noise measurement 
    Reply  Wed Jan 27 18:19:45 2016, gautam, Update, Green Locking, Lightwave frequency noise measurement 
       Reply  Thu Jan 28 00:29:30 2016, gautam, Update, Green Locking, Lightwave frequency noise measurement PLL_OLG.pdfFreq_noise_comparison.pdf
          Reply  Mon Feb 1 18:11:25 2016, gautam, Update, Green Locking, Lightwave frequency noise measurement Free_running_frequency_noise_comparison.pdf
       Reply  Fri Jan 29 14:43:47 2016, Steve, Update, Green Locking, Innolight laser is 10 years old inno1W.pdf
          Reply  Fri Jan 29 16:55:27 2016, rana, Update, Green Locking, Innolight laser is 10 years old 
             Reply  Sat Jan 30 00:12:22 2016, gautam, Update, Green Locking, Innolight laser is 10 years old RIN_comparison.pdf
                Reply  Sat Jan 30 09:56:24 2016, Koji, Update, Green Locking, Innolight laser is 10 years old 
                   Reply  Mon Feb 1 15:16:28 2016, gautam, Update, Green Locking, Innolight laser is 10 years old 
                      Reply  Mon Feb 1 15:43:18 2016, Koji, Update, Green Locking, Innolight laser is 10 years old 
                         Reply  Fri Feb 5 15:02:13 2016, gautam, Update, Green Locking, X-end NE cable 
          Reply  Fri Feb 12 11:10:49 2016, Steve, Update, Green Locking, Innolight laser is 10 years old inno1W.jpg
          Reply  Mon Mar 21 14:29:32 2016, Steve, Update, Calibration-Repair, 1W Innolight laser repair diagnoses 
             Reply  Thu Mar 24 07:56:09 2016, Steve, Update, Calibration-Repair, NO Noise Eater for 1W Innolight 
                Reply  Mon Apr 11 17:03:41 2016, Steve, Update, Calibration-Repair, 1W Innolight repair completed 
Message ID: 11956     Entry time: Thu Jan 28 00:29:30 2016     In reply to: 11953     Reply to this: 11969
Author: gautam 
Type: Update 
Category: Green Locking 
Subject: Lightwave frequency noise measurement 

Summary of the work done today:

Alignment and other work on PSL table

As mentioned in a previous elog, the beatnote amplitude I obtained was tiny - so I checked the alignment of the two beams onto the PD. I did this as follows:

  • Checked the alignment of the two beams on the recombination BS. Moved the steering mirror for the PSL beam until the two were aligned, as verified by eye using an IR card
  • Turn the steering mirror just before the fast focusing lens and thorlabs PD (kept the fork fixed, just loosened the screw on the post to do this) such that the far-field alignment of the two beams could be checked. I used the BS to tweak this alignment as necessary
  • Iterate the previous two steps till I was happy with the alignment
  • Return steering mirror before the PD to its original position, tweak alignment until DC level on the PD was maximized (as verified using an oscilloscope) 
  • Adjust the HWP just after the lightwave laser such that the power arriving at the PD from the PSL beam and the lightwave beam were approximately equal - verified by blocking each beam and checking change in the DC level

After doing all of this, I found a beatnote at ~-10dBm at a temperature of 45.3002 degrees on the Lightwave. The DC level was ~8V (~4V contribution from each beam). 

PLL and frequency nosie measurements:

Pretty much the same procedure as that described in this elog was followed for setting up the PLL and taking the measurements, except that this time, I used the two SR560s in a better way to measure the open loop TF of the PLL. This measurement suggested a UGF of ~ 10kHz, which seems reasonable to me. I turned the 11MHz marconi off because some extra peaks were showing up in the beat signal spectrum. I judged that the beatnote was not large enough to require the use of an attenuator between the PD and the mixer. I was able to lock the PLL easily enough, and I've attached spectra of the control signal (both uncalibrated and calibrated). To calibrate the spectrum, I did a quick check to determine the actuator gain of the spare Lightwave laser, by sweeping the fast PZT with a low frequency (0.5Hz) 1Vpp sine wave, and looking at the peak in the beat signal spectrum move on the network analyzer. This admittedly rough calibration suggests that the coefficient is ~5MHz/V, consistent with the other Lightwave. Eric suggested a more accurate way to do this would be to match up spectra taken using this method and by locking the PLL by actuating on the FM input of the Marconi - I didn't try this, but given the relatively large low-frequency drifts of the beatnote that I was seeing, and that the control signal was regularly hitting ~2V (i.e shifting the frequency by ~10MHz), I don't think this is viable with a low MHz/V coefficient on the Marconi, which we found is desirable as described here

Bottom line:

The spare Lightwave frequency noise seems comparable to the other two measurements (see attachment #2). If anything, it is a factor of a few worse, though this could be due to an error in the calibration? I'm also not sure why the shapes of the spectra from today's measurement differ qualitatively from those in elog 11929 above ~7kHz. 

 

Some random notes:

  • Do we want to do an AM/PM characterization of the spare Lightwave laser as well? It might be easier to do the PM measurement while we have this measurement setup working
  • Yesterday, I noticed some peaks in the spectrum of the PD output while only the PSL beam was incident on it, at ~35MHz and ~70 MHz. They were pretty small (~-50dBm), but still clearly discernible over the analyzer noise floor. It is unclear to me what the source of these peaks are.
Attachment 1: PLL_OLG.pdf  54 kB  | Hide | Hide all
PLL_OLG.pdf
Attachment 2: Freq_noise_comparison.pdf  95 kB  Uploaded Sat Jan 30 00:25:18 2016  | Hide | Hide all
Freq_noise_comparison.pdf
ELOG V3.1.3-